the median class of a frequency distribution in 125-145. the frequency and cumulative frequency of the class preceding to the median class are 20 and 22 respectively. Find the sum of frequencies , if the median is 137
 

Grade:11

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Median class = 125 – 145
Frequency of median class (f) = 20
Cumulative frequency of the class preceding to the median class (cf) = 22
Median = 137
Class size (h) = 145 – 125 = 20
Lower Limit of median class (l) = 125
Let the sum of the frequencies be n
Median=l+\frac{n/2-cf}{f}h
137 =125+(((n/2)-22)/20)*20
137=125+(n/2)-22
n/2=137-125+22
n/2=34
n=68